# ReX: Causal Discovery based on Machine Learning and Explainability techniques

### v1

@@ 3 Proposed method
The proposed method ReX approaches the causal discovery problem in a series of steps (see Fig. 1). First, we fine-tune and train a series of regressors to predict each variable $X_{i}$, for $i=1,\dots,p$, using the remaining $p-1$ variables in the dataset, denoted as $X_{\backslash i}=\{X_{1},\dots,X_{i-1},X_{i+1},\ldots,X_{p}\}$. Once the regressors are fine-tuned and trained, we compute Shapley values to estimate the contribution of each variable to the prediction, applying a bootstrapping mechanism, i.e., iteratively using different samples of the dataset. Through this repeated sampling, ReX identifies a robust set of features that can be considered the potential causes (i.e., $pa(X_{i})$) for each target variable $X_{i}$. As a final step, results from the different regressors are combined to obtain a plausible causal graph, which is reviewed to direct edges and remove eventual cycles. Next, we describe these steps in detail.
@@
Figure 1: Summary of the different steps performed by ReX.
@@ 3.1 Models training
The first step in deriving the causal model is to train a separate model to predict each variable in the dataset, with hyper-parameter optimization (HPO). Before model training, the data is normalized by removing the mean and scaling to unit variance. To ensure robustness, we employ two complementary regressors: a deep feed-forward neural network (DFN) and gradient boosting trees (GBT). DFNs are highly flexible models capable of approximating any function, leveraging their capacity as universal approximators, as demonstrated by [23]. GBTs, known for their effectiveness in handling tabular continuous data and offering more interpretable outputs as shown in [24], serve to validate and complement the results obtained from the DFN models, with the XGBoost implementation used in this study. Although the ReX framework allows for the integration of additional regressors, the results from DFN and GBT have proven to be sufficiently robust for our purposes.
A total of $p$ DFN and $p$ GBT models are trained, where the target variable is feature $X_{i}$, for $i=1,\ldots,p$, and the input data consists of the remaining variables $X_{\backslash i}$. The DFN models include an additional additive noise variable $\varepsilon\sim\textrm{Uniform}(0,1)$ as input, which is re-generated at each training forward step. Each DFN and GBT model is trained with mean squared error (MSE) as loss function. These models act therefore as functions $\hat{\mkern-2.0muf\mkern-3.0mu_{i}}$ that map inputs $X_{\backslash i}$ to the estimate of the corresponding $X_{i}$, i.e., $\hat{\mkern-2.0muf\mkern-3.0mu_{i}}(X_{\backslash i},\varepsilon_{i})=\hat{X_{i}}\approx X_{i}$.
@@ 3.2 Bootstrapping and Initial Causal Graph Construction
In this phase, we implement the bootstrapping approach to construct an initial undirected causal graph. This process involves three main steps: computing SHAP values to assess feature importance, selecting potential parent features based on their impact, and updating an adjacency matrix that represents candidate causal connections. These steps are iteratively repeated in the bootstrapping loop (Algorithm 1) to build a stable causal structure.
@@
0:  Let $\hat{f}$ be a regressor model, fine tuned –DFN or GBT.
1:  Initialize an empty adjacency matrix $\mathbf{A}\in\mathbb{R}^{p\times p}$.
2:  for $t=1$ to $T$ (nr. of bootstrapping iterations) do
3:     Sample a subset of data $\mathbf{X}^{(t)}$ from $\mathbf{X}$ (observational data).
4:     for each feature $X_{i}$ do
5:        Fit SHAP explainer with $\mathbf{X}^{(t)}_{\smallsetminus i}$ and model $\hat{f}$, to obtain  $\Phi^{(i)}=\{\phi_{1}\ldots\phi_{i-1},\phi_{i+1}\ldots\phi_{p}\}$
6:        Run Algorithm 2 with $\Phi^{(i)}$ to obtain $\textit{pa}(X_{i})$
7:        for each pair of adjacent features $(X_{i},X_{j}\in\textit{pa}(X_{i}))$ do
8:           Increment the corresponding entry in $\mathbf{A}$:  $a_{i,j}\leftarrow a_{i,j}+1$
9:        end for
@@
12:  Normalize A by $T$: $a_{i,j}\leftarrow\tfrac{a_{i,j}}{T}$
13:  Filter edges in $\mathbf{A}$: $a_{i,j}\leftarrow 0$ if $a_{i,j}<\tau$
14:  return  Adjacency matrix $\mathbf{A}$ representing the causal graph
Algorithm 2 Parents selection from SHAP values
0:  Input array $\Phi=\{\phi_{1}\ldots\phi_{p-1}\}$, SHAP values
1:  $\Delta\leftarrow$ Pairwise Euclidean distances in $\Phi$
2:  $\Delta^{\prime}\leftarrow$ Sort diffs in $\Delta$ in descending order
3:  Initialize $\varepsilon$ to $\max(\Delta)+\text{constant}$
4:  repeat
5:     Run DBSCAN on $\Phi$ with parameter $\varepsilon$
6:     $n\leftarrow$ Number of clusters
7:     if $n=1$ then
8:        $\varepsilon\leftarrow\varepsilon-\max(\Delta^{\prime})$
9:        Remove $\max(\Delta^{\prime})$ from $\Delta^{\prime}$
10:     end if
11:  until $n>1$ or $\Delta^{\prime}$ = $\emptyset$
12:  if No clusters formed then
13:     return  None
14:  else
15:     return  Features in cluster with highest mean SHAP value
16:  end if
@@ 3.2.2 Parents Selection
In the second step of Algorithm 1, we apply the DBSCAN clustering algorithm to group features based on their SHAP values and select those that have a significant impact on predicting the target variable. Feature clusters are adjusted dynamically by decreasing the neighborhood radius parameter $\varepsilon$ until more than one group are detected, helping to separate influential features from less impactful ones. This clustering process effectively selects potential parent features, as it isolates variables that likely have causal influence on the target. Algorithm 2 outlines this clustering procedure in detail, and Appendix A describes the process with a visual example and additional details.
@@ 3.2.3 Adjacency Matrix Update and Graph Construction
With the influential features identified, we proceed to construct an initial causal graph by updating an adjacency matrix, which records candidate edges between features. In each bootstrapping iteration, a subset of the data $\mathbf{X}^{(t)}$ is sampled, and SHAP-based parent selection is performed. For each pair of selected features, the corresponding entry in the adjacency matrix $\mathbf{A}$ is incremented, capturing the frequency of selection across iterations. After bootstrapping, we apply a threshold $\tau$ to filter out edges with low frequencies, retaining only stable and meaningful connections. This thresholding process yields a plausible undirected causal graph, representing the initial structure without directed edges.
The percentage of rows to be selected (step 3 in Algorithm 1) is set to be $\geq 1-q^{1/r}$, where $q$ is the maximum acceptable probability that any given row is not selected in any iteration (typically $q=0.01$), and $r$ is the number of iterations defined for the bootstrapping mechanism (typically $r=100$).
@@ 3.3 Directing edges
The direction of each edge in the graph is established based on the method proposed by the Additive Noise Model (ANM) for structure identification as described in [3] and improved in [26]. This procedure is applied to each unoriented edge by fitting a regression model (either a Gaussian Process Regressor (GPR) or a Generalized Additive Model (GAM)) to approximate the dependent variable ($X_{j}$) as a function of the predictor ($X_{i}$) plus Gaussian noise. This process is repeated in the reverse direction, with $X_{i}$ as the dependent variable and $X_{j}$ as the predictor, for each selected pair of features.
To determine whether the variables are conditionally independent, we use the Hilbert-Schmidt Independence Criterion (HSIC) [27], specifically testing for independence between the residuals of the regression and the predictor variable. We adopt the Gaussian kernel for the HSIC independence test, which uses the median distance as the bandwidth, guaranteeing consistency in the test [28]. If the independence hypothesis is accepted, we infer that the causal direction is from the predictor to the dependent variable.
@@ 3.4 Final DAG
[…]
The SHAP discrepancy is introduced as a measure to assess how well the Shapley values of a feature explain the variability in the target variable. It quantifies the extent to which the contributions of a feature $X_{j}$ (captured by its Shapley values $\phi_{j}$) align with the actual values of $X_{i}$ (see Appendix B for more details), computed as follows:
$\delta_{j}^{(i)}=1-R^{2}(X_{i},\phi_{j})=\frac{\sum_{k=1}^{m}(x_{k,i}-\varphi_{k,j})^{2}}{\sum_{k=1}^{m}(x_{k,i}-\overline{X_{i}})^{2}},$ (11)
where $x_{k,i}$ is the $k$-th observation of $X_{i}$, $\varphi_{k,j}$ is the corresponding Shapley value of $X_{j}$ for predicting $X_{i}$, $\overline{X_{i}}$ is the mean of $X_{i}$, and $m$ is the total number of samples.
@@ 3.5 Visual example
[…]
@@ 4 Results on synthetic data
We first consider synthetic data for the evaluation, as it allows to operate under controlled assumptions, where the causal structure is fully specified, and the data generation process is explicitly governed by known structural equations. Consequently, there are no unobserved confounders, fulfilling the causal sufficiency assumption, and the structural equations are stable under interventions. Additionally, the data respects the faithfulness assumption, ensuring that all observed conditional independencies align with the causal structure of the DAG.
@@
All computations were performed on a system equipped with an Apple M2 Pro processor and 32 GB of RAM, without the use of parallelization or GPU acceleration.
Table 1: Average and standard deviation number of edges
in the generated DAGs for $p=10,15,20,25$.
@@ 4.2 Experimental Results
Figure 3: Effect of increasing the number of variables in the input dataset with ReX in the following metrics: (a) F1 score; (b) SHD; (c) SID; and (d) computation time. Results are computed for the five considered families of synthetic data.
@@ 4.2.1 Results by synthetic dataset
When evaluating these metrics per synthetic dataset family (see Appendix E, Fig. 12), we do not observe significant differences. The results indicate that ReX performs robustly across a range of complexities in data structure, showing consistent trends across various settings.
@@ 4.2.2 Comparison with other methods
Next, we compare ReX against the compared methods (see also Appendix D for sample DAGs generated by each method). We observe that ReX consistently shows strong performance, particularly in F1 score (Fig. 4a), where it achieves higher values relative to most other methods. Notably, ReX exhibits a lower variance than the other methods, with values tightly concentrated around the median. This stability across metrics suggests that ReX is less sensitive to variations in the data, resulting in more consistent performance. As suggested by the obtained F1 scores, ReX stands out in aligning well with the true number of edges in the causal graph (Fig. 4b), indicating its effectiveness at avoiding both excessive edge predictions (false positives) and missing edges (false negatives). Compared to other methods, it exhibits a narrower interquartile range, implying lower variability in edge prediction across different datasets. This consistency suggests a robustness in ReX ’s performance, with fewer instances of significant deviations from the true edge count.
@@
Figure 5: Benchmark between ReX and compared methods for (a) SHD and (b) SID metrics.
@@
5 Results with Single-cell protein network data
We apply ReX to the well-studied single-cell protein-signaling network dataset, which was originally analyzed by [32] to assess the performance of structure learning methods against a largely known ground truth. This dataset, commonly referred to as Sachs, consists of $m=7466$ measurements of protein concentrations for $p=11$ proteins: Plc, Pkc, Pka, PIP2, PIP3, Raf, Mek, Erk, Akt, P38, and Jnk. For our analysis, we utilize the cleaned version of the Sachs dataset, as provided in [33], and used in [34].

### v2

@@ 3 Proposed method
The proposed method ReX approaches the causal discovery problem in a series of steps (see Fig. 1). First, we fine-tune and train a series of regressors to predict each variable $X_{i}$, for $i=1,\ldots,p$, using the remaining p1p-1 variables in the dataset, denoted as $X_{\backslash i}=\{X_{1},\ldots,X_{i-1},X_{i+1},\ldots,X_{p}\}$. Once the regressors are fine-tuned and trained, we compute Shapley values to estimate the contribution of each variable to the prediction, applying a bootstrapping mechanism, i.e., iteratively using different samples of the dataset. Through this repeated sampling, ReX identifies a robust set of features that can be considered the potential causes (i.e., $\text{Pa}(X_{i})$) for each target variable $X_{i}$. As a final step, results from the different regressors are combined to obtain a plausible causal graph, which is reviewed to direct edges and remove eventual cycles. Next, we describe these steps in detail. It is worth noting that while the overall structure of the ReX pipeline—training predictive models, assessing feature contributions, and constructing a graph—is a general approach, the core novelty of ReX lies in its specific instantiation using Shapley values for robust feature impact assessment and parent selection, guided by the theoretical connections discussed in Section 2.2, and its subsequent steps for graph refinement.
@@
Figure 1: Overview of the ReX workflow. The process begins with training regressors (Section 3.1), followed by a bootstrapping procedure (Algorithm 1) that includes data sampling, computation of SHAP values ($\phi$) to assess feature impact (Section 3.2.1), selection of candidate parents based on these values (Algorithm 2, Section 3.2.2), and updating the adjacency matrix (Section 3.2.3). Edges are then oriented (Section 3.3), and outputs from multiple regressors are combined and pruned to produce the final DAG (Section 3.4). While parts of the pipeline are general, ReX is specifically designed to leverage Shapley values for assessing feature contributions, as motivated in Section $2.$
@@ 3.1 Models training
The first step in deriving the causal model is to train a separate model to predict each variable in the dataset, with hyper-parameter optimization (HPO). Before model training, the data is normalized by removing the mean and scaling to unit variance. To ensure robustness, we employ two complementary regressors: a deep feed-forward neural network (DFN) and gradient boosting trees (GBT). DFNs are highly flexible models capable of approximating any function, leveraging their capacity as universal approximators, as demonstrated by [24]. GBTs, known for their effectiveness in handling tabular continuous data and offering more interpretable outputs as shown in [25], serve to validate and complement the results obtained from the DFN models, with the XGBoost implementation used in this study. Although the ReX framework allows for the integration of additional regressors, the results from DFN and GBT have proven to be sufficiently robust for our purposes (see Appendix D).
[…]
@@ 3.2 Bootstrapping and initial causal graph construction
In this phase, we implement the bootstrapping approach to construct an initial undirected causal graph ($\widehat{G}_{\text{undir}}$). This process involves three main steps: computing SHAP values to assess feature importance, selecting potential parent features based on their impact, and updating an adjacency matrix that represents candidate causal connections. These steps are iteratively repeated in the bootstrapping loop (Algorithm 1) to build a stable causal structure.
@@
1:Let $\hat{f}$ be a fine-tuned regressor model (DFN or GBT)
2:Initialize empty adjacency matrix $\mathbf{A}\in\mathbb{R}^{p\times p}$
3:for $t=1$ to $T$ do $\triangleright$ Number of bootstrap iterations
4:  Sample a subset $\mathbf{X}^{(t)}$ from $\mathbf{X}$ (observational data)
5:  for each feature $X_{i}$ do
6:   Fit SHAP explainer with $\mathbf{X}^{(t)}_{\smallsetminus i}$ and model $\hat{f}$ to obtain $\Phi^{(i)}=\{\phi_{1},\ldots,\phi_{i-1},\phi_{i+1},\ldots,\phi_{p}\}$
7:   Run Algorithm 2 with $\Phi^{(i)}$ to obtain $\textit{Pa}(X_{i})$
8:   for each $X_{j}\in\textit{Pa}(X_{i})$ do
9:     Increment the corresponding entry: $a_{i,j}\leftarrow a_{i,j}+1$
@@
12:end for
13:Normalize: $a_{i,j}\leftarrow\tfrac{a_{i,j}}{T}$ for all $(i,j)$
14:Filter edges: $a_{i,j}\leftarrow 0$ if $a_{i,j}<\tau$
15:return Adjacency matrix $\mathbf{A}$ representing a stable undirected graph $\widehat{G}_{\text{undir}}$
Algorithm 2 Parent Selection from SHAP Values
1:Input array $\Phi=\{\phi_{1},\ldots,\phi_{p-1}\}$ (SHAP values)
[…]
4:Initialize $\zeta\leftarrow\max(\Lambda)+\text{constant}$
5:repeat
6:  Run DBSCAN on $\Phi$ with parameter $\zeta$
7:  $n\leftarrow$ Number of clusters
8:  if $n=1$ then
9:   $\zeta\leftarrow\zeta-\max(\Lambda^{\prime})$
10:   Remove $\max(\Lambda^{\prime})$ from $\Lambda^{\prime}$
11:  end if
12:until $n>1$ or $\Lambda^{\prime}=\emptyset$
13:if no clusters formed then return None
14:elsereturn Features in cluster with highest mean SHAP value
15:end if
@@ 3.2.2 Parents selection
In the second step of Algorithm 1, we apply the DBSCAN clustering algorithm to group features based on their SHAP values and select those that have a significant impact on predicting the target variable. Feature clusters are adjusted dynamically by decreasing the neighborhood radius parameter $\zeta$ until more than one group are detected, helping to separate influential features from less impactful ones. This clustering process effectively selects potential parent features, as it isolates variables that likely have causal influence on the target. Algorithm 2 outlines this clustering procedure in detail, and Appendix B describes the process with a visual example and additional details.
DBSCAN is selected for clustering SHAP values because, unlike methods such as K-Means, it does not require specifying the number of clusters upfront—a crucial advantage given the data-dependent and unknown grouping of feature importances. Moreover, DBSCAN intrinsically distinguishes influential parent candidates (high SHAP values) from negligible-impact features by labeling low-density points as noise, a capability partition-based methods lack. Algorithm 2 further improves DBSCAN by adaptively setting the neighborhood radius parameter $\zeta$ according to the empirical distribution of SHAP values, avoiding arbitrary thresholds and providing a principled distinction between salient and negligible feature contributions.
@@ 3.2.3 Adjacency matrix update and graph construction
With the influential features identified, we proceed to construct an initial causal graph by updating an adjacency matrix, which records candidate edges between features. In each bootstrapping iteration, a subset of the data $\mathbf{X}^{(t)}$ is sampled, and SHAP-based parent selection is performed. For each pair of selected features, the corresponding entry in the adjacency matrix $\mathbf{A}$ is incremented, capturing the frequency of selection across iterations. After bootstrapping, we apply a threshold $\tau$ to filter out edges with low frequencies, retaining only stable and meaningful connections. This thresholding process yields a plausible undirected causal graph, representing the initial structure without directed edges ($\widehat{G}_{\text{undir}}$). Thus, the stability of an edge in this initial graph is directly determined by the parameter $\tau$; an edge is included if its frequency of selection during the bootstrapping procedure (Algorithm 1) meets or exceeds this threshold, signifying its robustness against variations in data subsamples.
Step 3 of Algorithm 1 involves creating a bootstrap sample $X^{(t)}$ by selecting a subset of the original data samples (i.e., rows) from the full dataset $X$. To ensure robust coverage of the dataset across the $T$ bootstrap iterations, the proportion $c$ of original data samples to be included in each such bootstrap sample $X^{(r)}$ is set using the formula $c\geq 1-q^{1/T}$. In this formula, $q$ is the maximum acceptable probability that any given original data sample (row) is not selected across any of the $T$ bootstrap iterations (typically $q=0.01$), and $T$ is the total number of iterations for the bootstrapping mechanism (typically $T=50$, as specified for Algorithm 1).
The undirected graph $\widehat{G}_{\text{undir}}$ produced at this stage is therefore stable under bootstrap perturbations; only these vetted edges are passed to the subsequent orientation module of Section 3.3.
@@ 3.3 Directing edges
Once a stable undirected graph $\widehat{G}_{\text{undir}}$ has been obtained—i.e., after the bootstrapping procedure of Section 3.2.3 has normalized the adjacency matrix and applied the frequency threshold $\tau$ in Algorithm 1 (lines 12–13) to retain only edges with selection probability $f_{ij}\geq\tau$—the next step is to orient the remaining candidate edges. The direction of each such pre-identified edge $(X_{i},X_{j})$ is established using principles from Additive Noise Models (ANMs), as described in [3]. This ANM-based orientation procedure involves fitting regression models for both potential directions (e.g., $X_{j}$ as a function of $X_{i}$, and $X_{i}$ as a function of $X_{j}$).
To determine the causal direction, we then use the Hilbert-Schmidt Independence Criterion (HSIC) [28] to test for independence between the residuals of each regression and the respective predictor variable. For a potential edge $X_{i}\rightarrow X_{j}$, if the HSIC test accepts the hypothesis that the residuals from regressing $X_{j}$ on $X_{i}$ are independent of $X_{i}$, we infer the causal direction $X_{i}\rightarrow X_{j}$. Conversely, if the HSIC test indicates dependence between the residuals and $X_{i}$ (i.e., the null hypothesis of independence is rejected, typically if the test $p$-value is below a significance threshold), this direction $X_{i}\rightarrow X_{j}$ is considered not supported by the ANM assumption. The process is repeated for the reverse direction. Thus, the HSIC test serves to ascertain whether the statistical dependence flagged by the Shapley values aligns with an identifiable ANM structure in one direction over the other.
@@ 3.4 Final DAG
The methodology presented here culminates in the generation of the final DAG, denoted as $\mathcal{G}_{\text{{R{e}X}}}$. This final DAG is produced by taking the union ($\cup$) of the two DAGs $\mathcal{G}_{\text{{DFN}}}$ and $\mathcal{G}_{\text{{GBT}}}$, which are generated by the described procedure when using DFN and GBT as the regressors, respectively (see Appendix D for a performance comparison among $\mathcal{G}_{\text{{R{e}X}}}$, $\mathcal{G}_{\text{{DFN}}}$ and $\mathcal{G}_{\text{{GBT}}}$).
The union ($\cup$) is chosen to leverage the complementary strengths of these diverse model types in identifying potential causal links, aiming for a more comprehensive discovery of true edges. This strategy is empirically supported by a higher overall F1-score compared to using intersection or individual regressors, as detailed in Appendix D. The implementation provides a configurable parameter that allows users to choose between computing the union or the intersection, depending on their specific requirements.
The final DAG $\mathcal{G}_{\text{{R{e}X}}}$ must be acyclic to represent valid causal relationships. Hence, in those cases where the new proposed DAG contains cycles or bidirectional edges between two nodes, as a result of the union of two DAGs, we employ a strategy based on the use of the SHAP discrepancy $\delta_{j}^{(i)}$ to make it acyclic.
The SHAP discrepancy is introduced as a measure to assess how well the Shapley values of a feature explain the variability in the target variable. It quantifies the extent to which the contributions of a feature $X_{j}$ (captured by its Shapley values $\phi_{j}$) align with the actual values of $X_{i}$ (see Appendix C for more details), computed as follows:
$\delta_{j}^{(i)}=1-R^{2}(X_{i},\phi_{j})=\frac{\sum_{k=1}^{m}(x_{k,i}-\phi_{k,j})^{2}}{\sum_{k=1}^{m}(x_{k,i}-\overline{X_{i}})^{2}},$ (12)
where $x_{k,i}$ is the $k$-th observation of $X_{i}$, $\phi_{k,j}$ is the corresponding Shapley value of $X_{j}$ for predicting $X_{i}$, $\overline{X_{i}}$ is the mean of $X_{i}$, and $m$ is the total number of samples.
@@ 3.5 Visual example
Figure 2: A visual summary of the intermediate graphs generated by ReX to obtain the final DAG $\mathcal{G}_{\text{{R{e}X}}}$. (a)-(b) represent the unoriented graphs for each of the two regressors after running the steps described in Sections 3.1, 3.2.1, 3.2.2 and 3.2.3. (c)-(d) are the result of establishing the direction of each edge presented in (a)-(b), respectively, following the approach described in Section 3.3. (e) is the initial DAG resulting from the union of DAGs (c)-(d). (f) is final DAG $\mathcal{G}_{\text{{R{e}X}}}$ resulting from removing cycles and/or bidirectional edged from (e), following the steps described in Section 3.4. The true DAG is also depicted for comparison. Green arrows represent correctly predicted edges, orange ones edges with the incorrect direction, and red ones edges that are missing in the true DAG.
@@ 4 Results on synthetic data
We first consider synthetic data for the evaluation, as it allows to operate under controlled assumptions, where the causal structure is fully specified, and the data generation process (see Appendix G) is explicitly governed by known structural equations. Consequently, in this synthetic setup, there are no unobserved confounders (i.e., variables that are common causes of two or more measured variables which could induce spurious associations), thereby fulfilling the causal sufficiency assumption (which posits that all such common causes are indeed part of the observed dataset). The structural equations are also stable under interventions, while all other structural equations remain unchanged (i.e., actively setting a variable’s value to observe its effects on downstream variables, as in $P\bigl(Y\mid\mathrm{do}(X\!=\!x)\bigr)$ [1]). Furthermore, the data respects the faithfulness assumption (formally defined in Section 2 and Definition 2), meaning all observed conditional independencies accurately reflect the d-separations in the true causal graph, with no accidental independencies due to parameter choices.
@@
It is important to note that while these default configurations do not typically incorporate an explicit internal bootstrapping loop for edge stability in the manner of ReX, our method’s bootstrapping procedure (Algorithm 1) is an integral component designed to ensure the robustness of parent sets derived from SHAP values of its machine learning regressors (DFNs and GBTs), whose explanations can exhibit variability with data resampling.
All computations were performed on a system equipped with an Apple M2-Pro processor and 32 GB of RAM, without the use of parallelization or GPU acceleration.
Table 1: Average and standard deviation number of edges in the generated DAGs for $p=10,15,20,25$.
@@ 4.2 Experimental results
Figure 3: Effect of increasing the number of variables in the input dataset with ReX in the following metrics: (a) F1 score; (b) SHD to true DAG; (c) SID to true DAG; and (d) computation time. Results are computed for the five considered families of synthetic data.
@@ 4.2.1 Results by synthetic dataset
When evaluating these metrics per synthetic dataset family (see Appendix F, Fig. 14), we do not observe significant differences. The results indicate that ReX performs robustly across a range of complexities in data structure, showing consistent trends across various settings.
@@ 4.2.2 Comparison with other methods
Next, we compare ReX against the compared methods (see also Appendix E for sample DAGs generated by each method). We observe that ReX consistently shows strong performance, particularly in F1 score (Fig. 4a), where it achieves higher values relative to most other methods. Notably, ReX exhibits a lower variance than the other methods, with values tightly concentrated around the median. This stability across metrics suggests that ReX is less sensitive to variations in the data, resulting in more consistent performance. As suggested by the obtained F1 scores, ReX stands out in aligning well with the true number of edges in the causal graph (Fig. 4b), indicating its effectiveness at avoiding both excessive edge predictions (false positives) and missing edges (false negatives). Compared to other methods, it exhibits a narrower interquartile range, implying lower variability in edge prediction across different datasets. This consistency suggests a robustness in ReX ’s performance, with fewer instances of significant deviations from the true edge count. The wider spread in edge difference for some of the compared methods observed in Figures 4 and 5 reflect their operation under default parameterizations (as stated in Section 4). These default settings might not be universally optimized for sparsity across all diverse synthetic data generation processes we tested, potentially leading them to predict denser graphs in certain scenarios compared to ReX’s intrinsic mechanisms for controlling graph density.
@@
Figure 5: Benchmark between ReX and compared methods for (a) SHD to true DAG and (b) SID to true DAG metrics.
@@
5 Evaluation on biological and financial complex systems
5.1 Single-cell protein network data
To evaluate ReX’s performance on a real-world biological system and facilitate comparison within the broader causal discovery literature, we selected the Sachs et al. (2005) single-cell protein-signaling network dataset [31]. This dataset is a prominent benchmark due to several factors: it represents a complex, genuine biological system; it possesses a widely accepted, albeit simplified, ground truth causal network derived from experimental interventions, which is crucial for quantitative validation; and its inherent characteristics, such as potential non-linear interactions, make it a relevant testbed for methods like ReX designed to handle such complexities. For our analysis, we utilize the cleaned version of the Sachs dataset, as provided in [32], and used in [33].